OpenStudy (anonymous):

Prove that if n^2 - (n - 2)^2 is not divisible by 8 then n is even

OpenStudy (anonymous):

n^2 - (n - 2)^2=n^2-(n^2-2n+4)=n^2-n^2+2n-4=2n-4 then we have 2n-4 when n is even 2(2m)-4=4(m-1) hmm not convinced of the question , when m is odd then n^2 - (n - 2)^2 divisible by 8 then

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

n^2 - (n - 2)^2=n^2-(n^2-4n+4)=n^2-n^2+4n-4=4n-4 = 4(n-1) @Marki

OpenStudy (anonymous):

oh yes !

OpenStudy (anonymous):

Thanks very much both :)